a) Kareena invested an amount of money at an interest rate of 4.5% per year. After one year, she received 117 dollars interest. Work out the amount of money kareena invested.
b) Ravi invested an amount of money at an interest rate of 4% per year. At the end of one year, interest was added to his account and the total amount in his account was 3328 dollars. Work out the amount of money Ravi invested.
: a) Kareena invested an amount of money at an interest rate of 4.5% per year. After one year, she received 117 dollars interest. Work out the amount of money kareena invested.
b) Ravi invested an amount of money at an interest rate of 4% per year. At the end of one year, interest was added to his account and the total amount in his account was 3328 dollars. Work out the amount of money Ravi invested.

Answer by vleith(1172)

Interest formula is given by
FV = PV(1+r)^n
You are asked to find original investment, given the interest for a single year.
Interest = PV(r)
117 = PV(0.045)
117/0.045 = PV
2600 = PV

2)
FV = PV(1+r)^n
3328 = PV(1+0.04)^1
3328/1.04 = PV
3200 = PV